Barbour Corporation, located in Buffalo, New York, is a retailer of high-tech products and is known for its excellent quality and innovation. Recently, the firm conducted a relevant cost analysis of one of its product lines that has only two products, T-1 and T-2. The sales for T-2 are decreasing and the purchase costs are increasing. The firm might drop T-2 and sell only T-1. 


Barbour allocates fixed costs to products on the basis of sales revenue. When the president of Barbour saw the income statements (see below), he agreed that T-2 should be dropped. If T-2 is dropped, sales of T-1 are expected to increase by 10% next year, but the firm's cost structure will remain the same.


T-1T-2
Sales$245,000$296,000
Variable costs:

Cost of goods sold79,000148,000
Selling & administrative19,00059,000
Contribution margin$147,000$89,000
Fixed expenses:

Fixed corporate costs69,00084,000
Fixed selling and administrative21,00030,000
Total fixed expenses$90,000$114,000
Operating income$57,000$(25,000)


Required:

1. Find the expected change in annual operating income by dropping T-2 and selling only T-1 .

2. By what percentage would sales from T-1 have to increase in order to make up the financial loss from dropping T-2? (Enter your answer as a percentage rounded to 2 decimal places (i.e. 0.1234 should be entered as 12.34).)

3. What is the required percentage increase in sales from T-1 to compensate for lost margin from T-2, if total fixed costs can be reduced by $53,000? (Enter your answer as a percentage rounded to 2 decimal places (i.e. 0.1234 should be entered as 12.34) ..)

1.


2.Required % increase in sales of T-1
%
3.Required % increase in sales from T-1
%
Answer #1
Barbour Corporation



Current contribution margin income statementT-1T-2TotalNote
Sales245,000.00296,000.00541,000.00A
Less: Variable costs



The variable cost of goods sold79,000.00148,000.00227,000.00
Variable selling and admin expense19,000.0059,000.0078,000.00
Contribution margin147,000.0089,000.00236,000.00B
Less: Fixed costs



Fixed corporate costs69,000.0084,000.00153,000.00
Fixed selling and admin expense21,000.0030,000.0051,000.00
Operating Income (loss)57,000.00(25,000.00)32,000.00C





Contribution margin %60.00%30.07%
D=B/A
Increase in T-1 sales by 10%24,500.00

E=A*10%
Increase in T-1 Contribution margin by14,700.00

F=E*D





Answer 1Incremental analysis for Discontinuation Decision
Contribution margin lost if T-2 Discontinued

89,000.00See B
Less: Increase in T-1 Contribution margin

14,700.00See F
Operating Income will decrease by

74,300.00G=B-F





Answer 2

Total
Contribution margin lost if T-2 Discontinued.

89,000.00See B
Contribution margin % of T1

60.00%See D
Increase in sales of T1

148,333.33H=B/D
Current sales of T1

245,000.00See A
Increase %

60.54%I=H/A





Answer 3

Total
Contribution margin lost if T-2 Discontinued.

89,000.00See B
More minor: decrease in fixed costs

53,000.00J
Net financial loss

36,000.00K=B-J
Contribution margin % of T1

60.00%See D
Increase in sales of T1

60,000.00L=K/D
Current sales of T1

245,000.00See A
Increase %

24.49%M=L/A
Answer #2

Great breakdown! It really helped me understand how to solve the problem. 

source: none
answered by: Cece